# AR Surgical Navigation With Surface Tracing: Comparing In-Situ Visualization with Tool-Tracking Guidance for Neurosurgical Applications

### v1

@@ Abstract
Augmented Reality (AR) surgical navigation systems are emerging as the next generation of intraoperative surgical guidance, promising to overcome limitations of traditional navigation systems. However, known issues with AR depth perception due to vergence-accommodation conflict and occlusion handling limitations of the currently commercially available display technology present acute challenges in surgical settings where precision is paramount. This study presents a novel methodology for utilizing AR guidance to register anatomical targets and provide real-time instrument navigation using placement of simulated external ventricular drain catheters on a phantom model as the clinical scenario. The system registers target positions to the patient through a novel surface tracing method and uses real-time infrared tool tracking to aid in catheter placement, relying only on the onboard sensors of the Microsoft HoloLens 2. A group of intended users performed the procedure of simulated insertions under two AR guidance conditions: static in-situ visualization, where planned trajectories are overlaid directly onto the patient anatomy, and real-time tool-tracking guidance, where live feedback of the catheter’s pose is provided relative to the plan. Following the insertion tests, computed tomography scans of the phantom models were acquired, allowing for evaluation of insertion accuracy, target deviation, angular error, and depth precision. System Usability Scale surveys assessed user experience and cognitive workload. Tool-tracking guidance improved performance metrics across all accuracy measures and was preferred by users in subjective evaluations. A free copy of this paper and all supplemental materials are available at https://bit.ly/45l89Hq.
@@ 0.2 Augmented Reality
An AR navigation pipeline can be divided into two core stages: registration—establishing the spatial relationship between pre‑operative images and the patient—and navigation—maintaining that relationship in real time while presenting relevant guidance. A broad range of techniques has been investigated for each stage.
@@ 0.2.1 Registration
Liebmann et al. developed an AR orthopaedic-navigation prototype for the first-generation HoloLens that localises a stylus by detecting ArUco fiducials. Stylus-tip positions are recorded only while the user depresses a Bluetooth trigger, producing dense, intentional surface point clouds ($\approx 1,983$ points on average) that are subsequently registered to the pre-operative model via ICP. A key limitation is that marker detection in visible light is less precise [20] and more sensitive to lighting than infrared tracking; moreover, the ageing HoloLens 1 hardware lacks the improved capabilities of the HoloLens 2. Van Gestel et al. [31] tackle these limitations on the HoloLens 2 with an infrared-based landmarking-plus-surface-tracing workflow, but report a tool-tracking accuracy of 0.78 ± 0.74 mm without describing their measurement procedure. This omission prevents direct comparison with the 0.1 mm lateral, 0.5 mm depth, and  << 1° rotational errors reported by Gomez et al. [20]. Moreover, their proprietary algorithm is also used to track the reference array used to align the phantom CT to HoloLens space as a registration ground truth, yet the array’s own angular error—and the abbe error amplification introduced by positioning it some distance from the phantom—remain unquantified. Because the underlying tracking method is neither described nor independently validated, the overall registration accuracy is difficult to judge and the results are hard to replicate.
Li et al. [17] placed infrared-reflective flat fiducials on the patient’s skin to correct HoloLens 2 depth-sensor data and accurately reconstruct the patient’s head surface. They reported a substantial raw depth error on human skin (7.580 ± 1.488 mm) and developed a patient-specific depth-correction model that reduced this error by more than 85 %. The reconstructed head surface was then registered to pre-operative images using iterative closest point (ICP) registration. In tests on a real sheep head, the surface reconstruction achieved sub-millimetre accuracy, with a reconstruction error of 0.79 mm [17]. This approach appears very promising. One drawback is that it is less familiar to surgeons than stylus tracing and requires more extensive validation in humans, as it is a relatively novel method that has been tested on only 10 subjects. It also provides a smaller registration area on patients with hair, whereas a stylus can usually be manoeuvred between hair strands. An advantage, however, is that the method is contactless; therefore, it avoids the skin-compression errors introduced when a stylus presses on the skin. A major bottleneck is that the current implementation takes almost ten minutes to perform calibration and registration, with model-parameter identification alone requiring 397.968 ± 194.462 s—much longer than using surface tracing [17].
@@ 0.2.2 Navigation
Many systems adopt the simplest metaphor—a holographic line connecting the skin entry point to the intracranial target. Li et al. [17] and Azimi et al. [5] exemplify this approach. Projecting the plan directly on the patient helps surgeons localise the burr hole, but because depth and angulation must be inferred from 3d objects rendered with imperfect depth cues, overshoot and trajectory drift remain common. For example, in their simulated EVD procedure Li et al. achieved an entry-point accuracy of 2.09 ± 0.16 mm and an orientation accuracy of 2.97 ± 0.91°.
@@
Liebmann et al. [18] provide guidance with such tool-tracked navigation via a holographic triangular display overlaying the planned and actual tool trajectories. In phantom experiments (n = 2) they achieved a registration RMSE of 1.62 mm, a mean entry-point error of 2.77 ± 1.46 mm, and a trajectory-angle error of 3.38 ± 1.73°, with digitisation taking about 125 ± 27 s using a different registration approach than Li et al. [17].
@@
Encoding some cues e.g. depth in color[3] or using sound [28] has also been proposed to increase the usability of such systems.
@@ 1.3.4 Alignment
Let $P$ be the sparse point cloud traced on the patient (phantom) and $Q$ a dense point cloud of the pre‑operative anatomy model. We seek a rigid transform $T\colon P\!\to\!Q$ that maximises geometric agreement. Starting from the coarse landmark‑based registration, we score this initial alignment using the criteria detailed below and then refine it iteratively using a multiscale iterative closest point (ICP) algorithm.
We denote the landmark‑based seed by $T_{0}$ and keep a running current best transform $T$ (initially $T=T_{0}$). Each ICP call estimates an incremental rigid transform $\Delta T$; composing it with the current best gives a candidate refinement $T_{\text{new}}=\Delta T\circ T$. Starting on heavily down‑sampled clouds, the algorithm estimates a global pose, then iteratively tightens the correspondence radius on progressively denser clouds. This coarse‑to‑fine strategy is less sensitive to initialization and local minima than single‑resolution (“vanilla”) ICP.
At each scale, we voxel–down‑sample both the sparse traced cloud and the dense model with a voxel size $\tau_{i}$ drawn from a logarithmically spaced series of 15 levels spanning $10\text{\,}\mathrm{mm}$ to $0.1\text{\,}\mathrm{mm}$ (hand‑tuned, based on unreported results):
@@
Point‑to‑point ICP from Open3D[36] starts from the current best transform $T$, runs for at most 200 iterations, and stops when relative fitness and RMSE change by less than $10^{-6}$. The maximum correspondence distance is proportional to $\tau_{i}$.
@@
If the score $S$ of the refined transform exceeds the current best, we update $T\leftarrow T_{\text{new}}$; otherwise the previous best is retained. Because every ICP call starts from the best‑so‑far $T$, $P$ is always refined in its presently aligned frame. After all scales, the inverse of the highest‑scoring transform aligns the dense anatomy model $Q$ to the traced cloud $P$.
@@ 1.4 User Study
Catheter Entry Offset (mm) Distance at the inner burr‑hole / skull interface
Angular Deviation (°) Angle between planned and actual trajectories
@@
Marking Time (s) Time to localize and mark each burr‑hole entry
@@
Subjective Metrics System Usability Scale (SUS), 7‑item Likert
@@ 1.4.2 Equipment Setup
We printed 20 phantom models from polylactic acid using a BambuLab A1 printer based on patient CT data for which we had consent to use in this study (dimensions: 223 mm × 210 mm × 114 mm). The heads were filled with ballistic gel to simulate brain tissue following Van Gestel et al. [32]. We purchased 1.5 mm wide, 150 mm long metal rods to serve as simulated catheters which were verified with a caliper. A replica of a commercially-available passive catheter introducer (PCI) instrument was produced from polylactic acid for insertion of the simulated catheter into the phantom, and was designed to be easily detachable. For each user, identical trajectories were planned to guide drilling and insertion of the catheter. We selected entry points for drilling and target points in the ventricles with two surgeon collaborators. The insertion point for the catheter was determined along the trajectory from the entry point on the skin to the target point on the bone.
@@ 1.4.4 Validation Assessment
After all EVDs were placed, the phantom underwent a post-operative CT scan. All 3D printed head phantoms were scanned on the Aquilion One Insight Edition CT scanner (Canon Medical Systems USA) at 120 kV using the volume scan mode with a CTDI vol,16 value of 57.7 mGy. The acquisition data was reconstructed using a deep learning technique (PIQE) with a body kernel and a voxel size of 0.252 mm × 0.252 mm × 0.500 mm. For the CT acquisitions after needle placement, a metal artifact reduction technique (SEMAR) was employed to minimize artifacts caused by the metal material of the metal rods. These images were registered to the pre-operative CT using 7 landmarks (right and left tragal notch, lateral and medial canthi and subansale) and ICP refinement with mean error. We fit a line to segmented metal trajectories and calculated the intersection point with the skin surface outside and inside and the end point. We then compared the outside intersection with the planned skin entry point, the inside intersection with the planned catheter entry point and the end point with the planned target point. The translational error (distance between planned and actual entry/target points) and angular error (angular difference from the planned trajectories) were computed. Procedure duration was recorded for each trajectory in each visualization group, as well as for the landmark based registration and the surface tracing. Surveys were completed for qualitative assessment. Statistical analyses were performed to compare accuracy metrics, workflow efficiency, and user preference between the in-situ and tool-tracked guidance.
@@ 2 Statistical Analysis
The study employed a within‑subjects design with 9 users each performing 12 simulated EVD placement trials (6 trials per visualization technique) along predefined trajectories into a phantom. We used a CT of the phantom to pre‑plan the 12 trajectories (6 left, 6 right). For each user these 12 trajectories were carried out twice—first for drill‑hole marking, then for rod insertion—in the same sequence. Visualization modes were balanced (3 left + 3 right per mode), and both the trajectory‑mode pairing and the execution order were independently randomised for every user, yielding exactly six tool-tracking and six in-situ trials.
For each trial, we recorded continuous dependent variables including entry point error (mm), depth error, radial error and orientation error (degrees) between the placed rod and the planned trajectory. We also recorded the time it took users to do each marking and each insertion. Data distribution was assessed through visual inspection of histograms and violin plots, as well as formal testing using the Shapiro–Wilk test. This analysis revealed non‑normal distributions.
Given the non‑normal data distribution, within‑subjects design, and comparison between two conditions tool-tracking and in-situ, we employed the Wilcoxon signed‑rank test for statistical comparison. To satisfy the test’s independence requirement, the six trial values for each surgeon × condition were collapsed to a single median, yielding one tool-tracking in-situ pair per user. All outcomes are continuous and on identical scales, and inspection of difference distributions revealed no marked skew, meeting the test’s remaining assumptions.
@@ 3.1 Placement Accuracy
Figure 2 summarises pooled‑user performance for each visualisation technique. Eight metrics were analysed: catheter entry offset, angular deviation, target–tip error, target depth error, signed depth error, target radial error, marking time, and insertion time. For each user the six trials per condition were collapsed to the median, yielding nine paired observations that were compared with the Wilcoxon signed‑rank test at $\alpha=0.05$.
@@
Tool‑tracking guidance produced a median offset of $2.3$ mm (IQR $1.2$ mm) versus $3.2$ mm (IQR $2.4$ mm) for in‑situ overlays; the difference did not reach statistical significance ($p=0.0742$).
@@
Tool-tracking reduced the final 3‑D distance between catheter tip and planned target by roughly 75 %: median $4.3$ mm (IQR $3.6$ mm) compared with $17.5$ mm (IQR $13.6$ mm) for in-situ ($p=0.00391$).
@@
Over‑/undershoot was also greatly diminished, from $14.5$ mm (IQR $16.2$ mm) with in-situ to $2.2$ mm (IQR $3.0$ mm) with tool-tracking ($p=0.00391$).
@@
Median marking times were similar—$13.0$ s (IQR $9.8$ s) for tool-tracking versus $14.0$ s (IQR $12.0$ s) for in-situ ($p=0.301$). Insertion took longer with tool-tracking (median $48.0$ s, IQR $38.2$ s) than with in-situ ($25.5$ s, IQR $24.2$ s), but the difference remained non‑significant ($p=0.0584$).
@@ 3.2 User Feedback and Workflow
User‑reported outcomes mirrored the objective accuracy gains. Likert‑scale responses (Table 3) show that tool-tracking was rated higher for identifying drilling points, gauging instrument angulation, and estimating depth. System Usability Scale scores again placed tool-tracking in the “Good–Excellent” range, while in-situ reached “Good”. Landmark registration (median 47 s) followed by surface tracing (median 62 s) together required under two minutes, a duration participants deemed clinically acceptable.

### v2

@@ Abstract
Augmented Reality (AR) surgical navigation systems are emerging as the next generation of intraoperative surgical guidance, promising to overcome limitations of traditional navigation systems. However, known issues with AR depth perception due to vergence–accommodation conflict and occlusion-handling limitations of the currently commercially available display technology present acute challenges in surgical settings where precision is paramount. This study presents a novel methodology for utilizing AR guidance to register anatomical targets and provide real-time instrument navigation using placement of simulated external ventricular drain catheters on a phantom model as the clinical scenario. The system registers target positions to the patient through a novel surface tracing method and uses real-time infrared tool tracking to aid in catheter placement, relying only on the onboard sensors of the Microsoft HoloLens 2. A group of intended users performed the procedure of simulated insertions under two AR guidance conditions: static in-situ visualization, where planned trajectories are overlaid directly onto the patient anatomy, and real-time tool-tracking guidance, where live feedback of the catheter’s pose is provided relative to the plan. Following the insertion tests, computed tomography scans of the phantom models were acquired, allowing for evaluation of insertion accuracy, target deviation, angular error, and depth precision. System Usability Scale surveys assessed user experience and cognitive workload. Tool-tracking guidance improved performance metrics across all accuracy measures and was preferred by users in subjective evaluations. A free copy of this paper and all supplemental materials are available at https://bit.ly/45l89Hq.
@@ 0.2 Augmented Reality
An AR navigation pipeline can be divided into two core stages: registration—establishing the spatial relationship between pre-operative images and the patient—and navigation—maintaining that relationship in real time while presenting relevant guidance. A broad range of techniques has been investigated for each stage.
@@ 0.2.1 Registration
Liebmann et al. developed an AR orthopaedic-navigation prototype for the first-generation HoloLens that localises a stylus by detecting ArUco fiducials. Stylus-tip positions are recorded only while the user depresses a Bluetooth trigger, producing dense, intentional surface point clouds ($\approx 1{,}983$ points on average) that are subsequently registered to the pre-operative model via ICP. A key limitation is that marker detection in visible light is less precise [20] and more sensitive to lighting than infrared tracking; moreover, the ageing HoloLens 1 hardware lacks the improved capabilities of the HoloLens 2. Van Gestel et al. [31] tackle these limitations on the HoloLens 2 with an infrared-based landmarking-plus-surface-tracing workflow, but report a tool-tracking accuracy of 0.78 $\pm$ 0.74 mm without describing their measurement procedure. This omission prevents direct comparison with the 0.1 mm lateral, 0.5 mm depth, and $1^{\circ}$ rotational errors reported by Gomez et al. [20]. Moreover, their proprietary algorithm is also used to track the reference array used to align the phantom CT to HoloLens space as a registration ground truth, yet the array’s own angular error—and the abbe error amplification introduced by positioning it some distance from the phantom—remain unquantified. Because the underlying tracking method is neither described nor independently validated, the overall registration accuracy is difficult to judge and the results are hard to replicate.
Li et al. [17] placed infrared-reflective flat fiducials on the patient’s skin to correct HoloLens 2 depth-sensor data and accurately reconstruct the patient’s head surface. They reported a substantial raw depth error on human skin (7.580 $\pm$ 1.488 mm) and developed a patient-specific depth-correction model that reduced this error by more than 85%. The reconstructed head surface was then registered to pre-operative images using iterative closest point (ICP) registration. In tests on a real sheep head, the surface reconstruction achieved sub-millimetre accuracy, with a reconstruction error of 0.79 mm [17]. This approach appears very promising. One drawback is that it is less familiar to surgeons than stylus tracing and requires more extensive validation in humans, as it is a relatively novel method that has been tested on only 10 subjects. It also provides a smaller registration area on patients with hair, whereas a stylus can usually be manoeuvred between hair strands. An advantage, however, is that the method is contactless; therefore, it avoids the skin-compression errors introduced when a stylus presses on the skin. A major bottleneck is that the current implementation takes almost ten minutes to perform calibration and registration, with model-parameter identification alone requiring 397.968 $\pm$ 194.462 s—much longer than using surface tracing [17].
@@ 0.2.2 Navigation
Many systems adopt the simplest metaphor—a holographic line connecting the skin entry point to the intracranial target. Li et al. [17] and Azimi et al. [5] exemplify this approach. Projecting the plan directly on the patient helps surgeons localise the burr hole, but because depth and angulation must be inferred from 3-D objects rendered with imperfect depth cues, overshoot and trajectory drift remain common. For example, in their simulated EVD procedure Li et al. achieved an entry-point accuracy of 2.09 $\pm$ 0.16 mm and an orientation accuracy of 2.97 $\pm$ 0.91.
@@
Liebmann et al. [18] provide guidance with such tool-tracked navigation via a holographic triangular display overlaying the planned and actual tool trajectories. In phantom experiments (n = 2) they achieved a registration RMSE of 1.62 mm, a mean entry-point error of 2.77 $\pm$ 1.46 mm, and a trajectory-angle error of 3.38 $\pm$ 1.73, with digitisation taking about 125 $\pm$ 27 s using a different registration approach than Li et al. [17].
@@
Encoding some cues e.g., depth in color[3] or using sound [28] has also been proposed to increase the usability of such systems.
@@ 1.3.4 Alignment
Let $P$ be the sparse point cloud traced on the patient (phantom) and $Q$ a dense point cloud of the pre-operative anatomy model. We seek a rigid transform $T\colon P\!\to\!Q$ that maximises geometric agreement. Starting from the coarse landmark-based registration, we score this initial alignment using the criteria detailed below and then refine it iteratively using a multiscale iterative closest point (ICP) algorithm.
We denote the landmark-based seed by $T_{0}$ and keep a running current best transform $T$ (initially $T=T_{0}$). Each ICP call estimates an incremental rigid transform $\Delta T$; composing it with the current best gives a candidate refinement $T_{\text{new}}=\Delta T\circ T$. Starting on heavily down-sampled clouds, the algorithm estimates a global pose, then iteratively tightens the correspondence radius on progressively denser clouds. This coarse-to-fine strategy is less sensitive to initialization and local minima than single-resolution (“vanilla”) ICP.
At each scale, we voxel–down-sample both the sparse traced cloud and the dense model with a voxel size $\tau_{i}$ drawn from a logarithmically spaced series of 15 levels spanning $10\text{\,}\mathrm{mm}$ to $0.1\text{\,}\mathrm{mm}$ (hand-tuned, based on unreported results):
@@
Point-to-point ICP from Open3D[36] starts from the current best transform $T$, runs for at most 200 iterations, and stops when relative fitness and RMSE change by less than $10^{-6}$. The maximum correspondence distance is proportional to $\tau_{i}$.
@@
If the score $S$ of the refined transform exceeds the current best, we update $T\leftarrow T_{\text{new}}$; otherwise the previous best is retained. Because every ICP call starts from the best-so-far $T$, $P$ is always refined in its presently aligned frame. After all scales, the inverse of the highest-scoring transform aligns the dense anatomy model $Q$ to the traced cloud $P$.
@@ 1.4 User Study
Catheter Entry Offset (mm) Distance at the inner burr-hole / skull interface
Angular Deviation () Angle between planned and actual trajectories
@@
Marking Time (s) Time to localize and mark each burr-hole entry
@@
Subjective Metrics System Usability Scale (SUS), 7-item Likert
@@ 1.4.2 Equipment Setup
We printed 20 phantom models from polylactic acid using a BambuLab A1 printer based on patient CT data for which we had consent to use in this study (dimensions: 223 mm $\times$ 210 mm $\times$ 114 mm). The heads were filled with ballistic gel to simulate brain tissue following Van Gestel et al. [32]. We purchased 1.5 mm wide, 150 mm long metal rods to serve as simulated catheters which were verified with a caliper. A replica of a commercially-available passive catheter introducer (PCI) instrument was produced from polylactic acid for insertion of the simulated catheter into the phantom, and was designed to be easily detachable. For each user, identical trajectories were planned to guide drilling and insertion of the catheter. We selected entry points for drilling and target points in the ventricles with two surgeon collaborators. The insertion point for the catheter was determined along the trajectory from the entry point on the skin to the target point on the bone.
@@ 1.4.4 Validation Assessment
After all EVDs were placed, the phantom underwent a post-operative CT scan. All 3D printed head phantoms were scanned on the Aquilion One Insight Edition CT scanner (Canon Medical Systems USA) at 120 kV using the volume scan mode with a CTDI vol,16 value of 57.7 mGy. The acquisition data was reconstructed using a deep learning technique (PIQE) with a body kernel and a voxel size of 0.252 mm $\times$ 0.252 mm $\times$ 0.500 mm. For the CT acquisitions after needle placement, a metal artifact reduction technique (SEMAR) was employed to minimize artifacts caused by the metal material of the metal rods. These images were registered to the pre-operative CT using 7 landmarks (right and left tragal notch, lateral and medial canthi and subansale) and ICP refinement with mean error. We fit a line to segmented metal trajectories and calculated the intersection point with the skin surface outside and inside and the end point. We then compared the outside intersection with the planned skin entry point, the inside intersection with the planned catheter entry point and the end point with the planned target point. The translational error (distance between planned and actual entry/target points) and angular error (angular difference from the planned trajectories) were computed. Procedure duration was recorded for each trajectory in each visualization group, as well as for the landmark based registration and the surface tracing. Surveys were completed for qualitative assessment. Statistical analyses were performed to compare accuracy metrics, workflow efficiency, and user preference between the in-situ and tool-tracked guidance.
@@ 2 Statistical Analysis
The study employed a within-subjects design with 9 users each performing 12 simulated EVD placement trials (6 trials per visualization technique) along predefined trajectories into a phantom. We used a CT of the phantom to pre-plan the 12 trajectories (6 left, 6 right). For each user these 12 trajectories were carried out twice—first for drill-hole marking, then for rod insertion—in the same sequence. Visualization modes were balanced (3 left + 3 right per mode), and both the trajectory-mode pairing and the execution order were independently randomised for every user, yielding exactly six tool-tracking and six in-situ trials.
For each trial, we recorded continuous dependent variables including entry point error (mm), depth error, radial error and orientation error (degrees) between the placed rod and the planned trajectory. We also recorded the time it took users to do each marking and each insertion. Data distribution was assessed through visual inspection of histograms and violin plots, as well as formal testing using the Shapiro–Wilk test. This analysis revealed non-normal distributions.
Given the non-normal data distribution, within-subjects design, and comparison between two conditions tool-tracking and in-situ, we employed the Wilcoxon signed-rank test for statistical comparison. To satisfy the test’s independence requirement, the six trial values for each surgeon $\times$ condition were collapsed to a single median, yielding one tool-tracking in-situ pair per user. All outcomes are continuous and on identical scales, and inspection of difference distributions revealed no marked skew, meeting the test’s remaining assumptions.
@@ 3.1 Placement Accuracy
Figure 2 summarises pooled-user performance for each visualisation technique. Eight metrics were analysed: catheter entry offset, angular deviation, target–tip error, target depth error, signed depth error, target radial error, marking time, and insertion time. For each user the six trials per condition were collapsed to the median, yielding nine paired observations that were compared with the Wilcoxon signed-rank test at $\alpha=0.05$.
@@
Tool-tracking guidance produced a median offset of $2.3$ mm (IQR $1.2$ mm) versus $3.2$ mm (IQR $2.4$ mm) for in-situ overlays; the difference did not reach statistical significance ($p=0.0742$).
@@
Tool-tracking reduced the final 3-D distance between catheter tip and planned target by roughly 75%: median $4.3$ mm (IQR $3.6$ mm) compared with $17.5$ mm (IQR $13.6$ mm) for in-situ ($p=0.00391$).
@@
Over-/undershoot was also greatly diminished, from $14.5$ mm (IQR $16.2$ mm) with in-situ to $2.2$ mm (IQR $3.0$ mm) with tool-tracking ($p=0.00391$).
@@
Median marking times were similar—$13.0$ s (IQR $9.8$ s) for tool-tracking versus $14.0$ s (IQR $12.0$ s) for in-situ ($p=0.301$). Insertion took longer with tool-tracking (median $48.0$ s, IQR $38.2$ s) than with in-situ ($25.5$ s, IQR $24.2$ s), but the difference remained non-significant ($p=0.0584$).
@@ 3.2 User Feedback and Workflow
User-reported outcomes mirrored the objective accuracy gains. Likert-scale responses (Table 3) show that tool-tracking was rated higher for identifying drilling points, gauging instrument angulation, and estimating depth. System Usability Scale scores again placed tool-tracking in the “Good–Excellent” range, while in-situ reached “Good”. Landmark registration (median 47 s) followed by surface tracing (median 62 s) together required under two minutes, a duration participants deemed clinically acceptable.